0.10 mol of argon gas is admitted to an evacuated 70 cm3 container at 20 ? C. The gas then undergoes an isothermal expansion to a volume of 200 cm3 . What is the final pressure of the gas? Express your answer using two significant figures.
Answer #1

Isothermal expansion means temperature is constant

Using ideal gas law:

PV = nRT

P2/P1 = V1/V2

P2 = P1*(V1/V2)

Now using given values:

P1 = n*R*T1/V1

V1 = 70 cm^3

V2 = 200 m^3

T1 = 20 C = 293 K

n = 0.10 mol

So,

P2 = (n*R*T1/V1)*(V1/V2)

P2 = n*R*T1/V2

So,

P2 = 0.10*8.314*293/(200*10^-6)

P2 = 1.218*10^6 Pa = 1.2*10^6 Pa

Please Upvote.